OpenStudy (anonymous):

in a study of 82 video game players, the researchers found that the ages of these players were normally distributed, with a mean age of 17 years and a standard deviation of 3 years. determine if their were 15 video game players in this study over the age of 20. justify your answer

OpenStudy (lilai3):

congrats on your first question

OpenStudy (anonymous):

Z = ( 20 - 17 ) / 3 = 1 P ( Z > 1 ) = .1587 => .1587 * 82 = 13 gamers Thus there are 13 gamers over the age of 20.